OpenStudy (anonymous):

Five years ago, Ana was twice as old as Jamie was then. Fifteen years from now, Jamie will be as old as Ana was ten years ago. How old is Jamie now?

OpenStudy (anonymous):

the answer is

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

12

OpenStudy (anonymous):

@Wherewillweland ???

OpenStudy (anonymous):

\[(A - 5) = 2 (J - 5), J + 15 = A - 10\]\[A=55, J=30\]

OpenStudy (anonymous):

how did you get 12 @smartguy1124 and thanks @robtobey

OpenStudy (anonymous):

@ Wherewillweland You're welcome.

OpenStudy (anonymous):

he right subtacted sorry